HIGH RESOLUTION JET COOLED CAVITY RINGDOWN SPECTROSCOPY OF THE Ã STATE 3_{0}^{1} BAND OF THE NO_{3} RADICAL

\underline{TERRANCE J. CODD}, MOURAD ROUDJANE and TERRY A. MILLER, Laser Spectroscopy Facility, The Ohio State University, Columbus, Ohio 43210.

The ^{2}E′′Ã state of NO_{3} is doubly degenerate and is therefore subject to Jahn-Teller (JT) distortion. In the Ã state there are two JT active modes, ν_{3} and ν_{4} (e′ stretch and in plane bend respectively). Theoretical work has predicted that the JT effect in the Ã state should be quite strong and approach the static case (D≥1) where the molecule is permanently distorted to a lower symmetry geometry.^{a}^{b}^{c} A moderate resolution spectrum of the Ã state showed a feature that we tentatively assigned as the 3_{0}^{1} band based on position and band contour.^{d} Using high resolution cavity ringdown spectroscopy we have now obtained a rotationally resolved spectrum of this band. The analysis of this band has been commenced using an oblate symmetric top Hamiltonian with spin-rotation terms. This analysis supports the assignment of this band to the a′′_{1} vibronic component of the 3_{0}^{1} band. So far, the spectrum shows no evidence of a large geometric distortion of the molecule. Some lines appear to be split, as was previously observed in the 4_{0}^{1} and 4_{0}^{2} bands,^{e} and the possible sources of this splitting are being investigated.

aJ.F. Stanton, 66^{th} OSU International Symposium on Molecular Spectroscopy, The Ohio State University, Columbus Ohio, 2011, TJ-03

bW. Eisfeld, K. Morokuma, J. Chem. Phys. 114, 9430 (2001)

cS. Faraji, H. Köppel, W. Eisfeld, S. Mahapatra, J. Chem. Phys. 347, 110 (2008)

dT.J. Codd, M.W. Chen, T.A. Miller, 66^{th} OSU International Symposium on Molecular Spectroscopy, The Ohio State University, Columbus Ohio, 2011, TD-06

eM.W. Chen, T.J. Codd, G. Just, T.A. Miller, OSU International Symposium on Molecular Spectroscopy, The Ohio State University, Columbus Ohio, 2011, TD-07